The company leases a machine which calls for a fixed charge plus a fee for each hour of usage. In November the company used the machine for 6,917 hours and paid $1,166. In December the company used the machine for 5,183 hours and paid $1,042. How much is the fixed monthly charge?
Answer #1

Monthly fixed charge = $671.36 or 671

Working

Cost No. of activities
A High Level $             1,166.00 6917
B Low Level $             1,042.00 5183
C=A-B Difference $                124.00                        1,734
A Cost difference $                124.00
B No. of activities difference 1734
C=A/B Variable cost per unit $                     0.07
High Level Low Level
A No. of activities 6917 5183
B Variable cost per unit $                0.0715 $                  0.0715
C=AxB Total Variable cost $                494.64 $                  370.64
D Total cost $             1,166.00 $              1,042.00
E=D-C Total fixed cost $                671.36 $                  671.36